OpenStudy (anonymous):

Solve the following system of equations: -7x+3y=-4 -8x+7y=24 x=? y=?

OpenStudy (anonymous):



OpenStudy (anonymous):

I found y, now just plug in the value of y, 8, into one of the original equations to find x.

OpenStudy (anonymous):

\[-7x+3y=-4\] \[-7x+3(8)=-4\] \[-7x+24=-4\] Can you continue from here? If you have any questions just ask away :D